Given the following function, find f (-4).
f(x) = 2−x^2 / x^3−x^2+6
f(-4)=

Answer by rfer(16322)

f(-4)=2--4^2/-4^3--4^2+6
f(-4)=2-16/-64-16+6
f(-4)=2+1/4-10
f(-4)=9/5-40/4
f(-4)=-31/4
f(-4)=-7 3/4